OpenStudy (anonymous):

WILL MEDAL!!! Zach borrowed four thousand dollars from his grandfather to buy a used car so he could travel to and from work. After working for 3 years, he has saved five thousand nine hundred twenty dollars to pay back to his grandfather. At what interest rate did Zach borrow the money? A. 11% B. 16% C. 23% D. 32%

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Very quick response Nono266 - please explain that to her. Thanks. Also, we do not provide answers without explanations on this site - violation of honor codes and they are watching for that. Happy studying

OpenStudy (anonymous):

I = p r t <=> r = I/ p * t The difference amount between borrowed and paid back is the interest: I = $5,920 - $4,000 = $1,920 => Interest rate: r = 1,920 / 4,000 * 3 = .16 = 16%

OpenStudy (anonymous):

sure thing @JuanitaM